OpenStudy (anonymous):

A sequence {an} is defined recursively, with a1 = -1, and, for n > 1, an = an-1 + (-1)n. Find the first five terms of the sequence. A) -1, 0, 1, 2, 3 B) -1, -2, -3, -4, -5 C) -1, 0, -1, 0, -1 D) -1, 1, -1, 1, -1

OpenStudy (johnweldon1993):

\(\large a_1 = -1\) and for \(\large n>1\) \(\large a_n = a_{n-1} + (-1)^n\) Like that?

OpenStudy (johnweldon1993):

I'm just seeing if I wrote it correctly as you have on your paper?

OpenStudy (anonymous):

yeah thats right

OpenStudy (johnweldon1993):

Okay So we know the first term is -1...we are given that...so \(\large a_1 = -1\) Now for \(\large a_2\) since we are now > 1 ...we use that equation \[\large a_n = a_{n-1} + -1^n\] So for n= 2 *the second term \[\large a_2 = a_{2-1} + -1^2\] \[\large a_2 = a_1 + -1^2\] We know \(\large a_1\) = -1 so we can replace that \[\large a_2 = -1 + -1^2\] What is -1^2? -1 * -1 = 1 right? so \[\large a_2 = -1 + 1 = ?\]